# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v2

@@ I Overview
A related phenomenon is the stochastic gravitational wave background induced by the enhanced curvature perturbation. In General Relativity, the curvature perturbation $\mathcal{R}$ and tensor perturbation $h_{ij}$ are only coupled at nonlinear order, and the tensor-scalar-scalar type $\sim h_{ij}\partial_{i}\mathcal{R}\partial_{j}\mathcal{R}$ can induce second-order GWs Matarrese et al. (1993, 1994, 1998); Noh and Hwang (2004); Carbone and Matarrese (2005); Nakamura (2007); Ananda et al. (2007); Osano et al. (2007); Baumann et al. (2007)  [^2] For GWs induced by other couplings like scalar-tensor-tensor or tensor-tensor-tensor, see for instance Refs. Chang et al. (2023); Yu and Wang (2023); Bari et al. (2024); Picard and Malik (2023).. Therefore, the curvature perturbation can source the GWs after it reenters the Hubble horizon, which is called scalar induced secondary GWs. The peak frequency, as well as the PBH mass generated by such a perturbation, are determined by the horizon scale at the reentry moment, which are connected by Saito and Yokoyama (2009)
@@
which is also used to describe the curvature perturbation on small scales for PBH formation Byrnes et al. (2012); Young and Byrnes (2013); Tada and Yokoyama (2015); Young and Byrnes (2015); Young et al. (2016); Franciolini et al. (2018); Ando et al. (2018a); Atal and Germani (2019); Passaglia et al. (2019); Özsoy and Tasinato (2023) as well as the induced GWs Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020). However, in many models, the non-linear parameter $f_{\mathrm{NL}}$ reaches $\mathcal{O}(1)$ Namjoo et al. (2013); Martin et al. (2013); Chen et al. (2013); Motohashi et al. (2015); Davies et al. (2022); Namjoo (2023); Namjoo and Nikbakht (2024), so the higher orders in (2) might be important, which are highly model-dependent. These higher order terms must be taken into account in the PBH formation. Fortunately, by $\delta N$ formalism, a fully nonlinear curvature perturbation can be derived for single field inflation and for the curvaton scenario, of which the PDF of $\mathcal{R}$ can be calculated analytically. Various methods are developed to calculate the PBH abundance for such a PDF, including the Press-Schechter-type formalism Biagetti et al. (2021); Gow et al. (2022); Ferrante et al. (2022) and the theory of peaks Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022). The Press-Schechter-type formalism for Gaussian curvature perturbation is thoroughly reviewed in Chapter 7 Young (2024), which is straightforward to be further extended to non-Gaussian case.
Considering non-Gaussianity in the PBH formation and induced GW generation is important in many theoretical issues and observational problems in cosmology.
@@ II.2.2 Threshold
We see that depending on whether $\mathscr{C}_{\ell}<4/3$ or $\mathscr{C}_{\ell}>4/3$, the maximum $\mathscr{C}(r_{m})$ corresponds to the maximum or minimum of $\mathscr{C}_{\ell}$. Only the former case, dubbed type I PBH Kopp et al. (2011), is consistent with the linear approximation. Type II PBHs with $\mathscr{C}_{\ell}>4/3$ are usually rarer, which we do not consider in this paper. Given the profile-dependent threshold on $\mathscr{C}$, the corresponding threshold on the linear compaction function $\mathscr{C}_{\ell}$ can be solved by (12)
@@ III.3 A general analysis on single-field inflation
$\mathcal{R}$ $=\delta N_{1}\left(\{\varphi,\pi\}\to\{\varphi_{t},\pi_{t}\}\right)+\delta N_{2}\left(\{\varphi_{t},\pi_{t}\}\to\{\varphi_{e},\pi_{e}\}\right)$ (44)
$\delta N_{1}$ $=\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi+\lambda_{\mp}\delta\varphi}{\pi+\lambda_{\mp}\varphi}\right)-\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda_{\mp}\varphi_{t}}\right)$ (45)
@@
Figure 4: The schematic potential (left panel) and the phase portrait (right panel) of the “smooth transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=-0.01$, $\varphi_{m}=0$. For fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$ (in the unit of $H$), respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Equal-$N$ lines are parallel to the fast-roll streams (i.e. $\pi\propto-3\varphi$) with almost equal intervals, which are not drawn for visual clearance. This clearly shows that the contribution to $\delta N$ from the ultra-slow-roll stage is negligible, and the contribution from the later stage preserved slow-roll-type Gaussianity.
@@
This result is easy to understand by looking at the phase portrait. For the “smooth transition” case, although the potential behaves differently for $\varphi<\varphi_{t}$ and $\varphi>\varphi_{t}$, but the non-attractor solution ($\pi=-3\varphi$) in the flat plateau is almost the same as the non-attractor solution in a slow-roll potential ($\pi=-\lambda_{+}^{\prime}\varphi\approx(3-\eta^{\prime})\varphi$). Therefore, this case is similar to the bumpy potential we studied in Section III.2, and (42) tells us $\mathcal{R}$ is nearly Gaussian with $f_{\mathrm{NL}}=-(6/5)\lambda_{-}^{\prime}$. The phase portrait Fig. 4 looks similar to that of a slow-roll inflation, as the ultra-slow-roll trajectories are similar to the fast-roll trajectories in an ordinary slow-roll inflation, which are parallel to the equal-$N$ lines. The difference in the $e$-folding number for two nearby trajectories are proportional to the difference in $\varphi$, therefore $\delta N\propto\delta\varphi$ which is a Gaussian relation. The well studied Starobinsky model Starobinsky (1992); Leach et al. (2001); Biagetti et al. (2018); Tasinato (2021); Pi and Wang (2023); Domènech et al. (2023) belongs to this limit.
Figure 5: The schematic potential (left panel) and the phase portrait (right panel) of the “sharp transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=0.01$, $\varphi_{m}=100$ (in the unit of $H$). For the fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$, respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Near the origin, the trajectories merge to a unique fast-roll trajectory with different $\pi$, which contribute negligibly to $\delta N$ in the subsequent evolution.
For the sharp transition case, the inflaton experiences an acceleration after the ultra-slow-roll stage, therefore the nearby trajectories, slightly shifted by the velocity difference $\delta\pi_{t}$ generated during the ultra-slow-roll stage, can also induce difference in $e$-folding number $\delta N$ in the later slow-roll stage. However, as $\pi_{t}\ll 1$, trajectories with different $\pi_{t}$’s near the origin are almost on the same fast-roll trajectory in the slow-roll era for $\varphi>\varphi_{t}$, as the trajectories near the $\varphi$-axis are perpendicular to it. On this trajectory, the difference in $N$ caused by $\delta\pi_{t}$ is negligible, and $\varphi=\varphi_{t}$ is approximately an equal-$N$ surface. Following the discussion in Section III.1, we will arrive at the non-Gaussian relation (30) straightforwardly.
@@ III.4 The curvaton scenario
where $r$ is roughly the cosmological parameter of the curvaton $\chi$ at its decay. Apparently such a nominal $f_{\mathrm{NL}}$ is not enough to calculate the PBH abundance, as $f_{\mathrm{NL}}$ is always larger than 1.
@@
The fully non-Gaussian curvature perturbation (49) seems complicated, but it has two interesting limiting cases for $r\ll 1$ and $r\to 1$, both of which have fruitful physical implications. In the former limit, the curvaton $\zeta$ is negligible at its decay, we have $r\ll 1$, and (49) can be expanded for small $r$:
@@
$\zeta=\frac{1}{3}\ln A=\frac{2}{3}\ln\left(1+\frac{\delta\chi}{\chi}\right).$ (54)
@@
$\zeta=\zeta_{\chi}(t,\mathbf{x})=\int^{\rho_{\chi}(t,\mathbf{x})}_{\rho_{\chi}(t)}\frac{\mathrm{d}\widetilde{\rho}_{\chi}}{3\widetilde{\rho}_{\chi}}=\frac{1}{3}\ln\frac{\rho_{\chi}(t,\mathbf{x})}{\rho_{\chi}(t)}=\frac{2}{3}\ln\left(1+\frac{\delta\chi}{\chi}\right),$ (55)
@@ IV PBH abundance with non-Gaussian curvature perturbation
where $\rho$ is the correlation between $X$ and $Y$, $j_{0}(z)=\sin z/z$ is the zeroth order spherical Bessel function, and we assume a spherical $\zeta$ profile of
@@
Another solution with $\mathscr{C}_{\ell}>4/3$ is discarded as it is for the Type II PBH. With (66), to change the variable from $\mathscr{C}_{\ell}$ to $M$, we have
@@
Apparently for $\lambda\to 0$, it degenerates to the Gaussian PDF. Roughly speaking, the variance of PDF (74) increases for a negative $\lambda$ (i.e. positive nominal $f_{\mathrm{NL}}$), which enhances PBH formation. As $\mathscr{C}_{\ell}$ increases, the PDF (74) finally approaches the Cauchy distribution $\mathbb{P}(\mathscr{C}_{\ell})\propto\mathscr{C}_{\ell}^{-2}$, which, however, is far from the Type-I-PBH formation region of $30.872<\mathscr{C}_{\ell}<4/$ 3 Biagetti et al. (2021). We calculated the PBH mass function $f_{\mathrm{PBH}}(M)$ for such a monochromatic power spectrum (71) in the ultra-slow-roll inflation ($\lambda=-3$), shown in the left panel of Fig. 6. We choose a typical PBH peak mass of $M_{k^{*}}=10^{-12}M_{\odot}$, and normalize the amplitude of the power spectrum $\mathcal{A_{R}}$ to be $4.151077\times 10^{-3}$, such that $f_{\mathrm{PBH}}=1$, i.e. PBHs can be all the dark matter. Interestingly, as is clearly seen in Fig. 6, for the narrow peak case, the critical infrared scaling $f_{\mathrm{PBH}}(M)\propto M^{1+1/\gamma}\sim M^{3.78}$ which originates from the universal critical behavior is almost parallel to the constraints from the 21-cm signal which is proportional to $M^{4}$ Mittal et al. (2022). So the entire asteroid-mass window is open for narrow-peak power spectrum even with non-Gaussianity.
@@ V Induced Gravitational Waves
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}\mathrm{g}}(vk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\lambda\int^{1+\lambda}_{|1-\lambda|}\mathrm{d}\rho~{}\frac{\mathcal{P}_{\mathcal{R}\mathrm{g}}(\lambda vk)\mathcal{P}_{\mathcal{R}\mathrm{g}}(\lambda vk)}{\lambda^{2}\rho^{2}}+\cdots\right)+\cdots.$ (83)
@@
Figure 7: Left: The spectrum of GWs induced by the enhanced narrow-peak power spectrum (71). The amplitude $\mathcal{A_{R}}$ for different parameters follows Table 1, such that PBHs $\sim 10^{-12}M_{\odot}$ are all the dark matter, shown in Fig.6. From right to left, the orange, magenta dashed, brown dot-dashed, cyan thick, blue dashed, purple dotted, black thick, gray, and black dotted curves are for $f_{\mathrm{NL}}=10^{3}$, $f_{\mathrm{NL}}=10^{2}$, $f_{\mathrm{NL}}=10$, ultra-slow-roll, $f_{\mathrm{NL}}=5/2$, $f_{\mathrm{NL}}=1$, Gaussian case ($f_{\mathrm{NL}}=0$), curvaton scenario with $r\to 1$, and $f_{\mathrm{NL}}=-5/4$, respectively. The curves are drown by Adshead et al. (2023), following Adshead et al. (2021), together with the designed power-law integrated sensitivity curve of space-borne interferometers LISA (gray thick) Bartolo et al. (2016), Taiji (gray dashed) Wang and Han (2021), TianQin (gray dotted) Liang et al. (2021), and DECIGO (gray thin) Schmitz (2021). Right: Two more realistic signals, the spectrum of GWs induced by ultra-slow-roll inflation with a sharp transition to slow-roll (blue), and a smooth transition to slow-roll, i.e. Starobinsky’s linear potential model (red). The right panel was drawn by (83) with the help of the SIGWfast Witkowski (2022) package.
@@
In the first expression, the numerical coefficient of the Gaussian contribution (“6.4”) comes from Ref. Pi and Sasaki (2020) with smoothing scale $\Delta=0.01$, while the coefficients of the $F_{\mathrm{NL}}^{2}$-term (“3.7”) and $F_{\mathrm{NL}}^{4}$-term (“3.9”) come from Ref. Adshead et al. (2021) (also with a width of $\Delta=0.01$). In the second equality, we use (79) for PBH as all the dark matter. From the above fitting formula, it is easy to see that $\Omega_{\mathrm{GW,peak}}h^{2}$ has a minimum of $\Omega_{\mathrm{GW,peak}}^{\text{(min)}}h^{2}\sim 8.4\times 10^{-11}$ when $f_{\mathrm{NL}}\sim 63$. After that, $\Omega_{\mathrm{GW}}$ only increase slightly, and finally approaches a constant of $1.1\times 10^{-10}$. This minimum is still well above the sensitivity curves of space-borne interferometers. Therefore, the detectability of scalar-induced GWs is robust against the nonlinear parameter $f_{\mathrm{NL}}$, $\lambda_{-}$ (see (34) and (42)), or $r$ (see (48) and (49)), which made it an important scientific goal of LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021) to probe the PBH dark matter.
@@
We should warn that the $\delta$-function peak we considered above is not physical. For more realistic models, the enhancement and decay of the power spectrum around the peak is milder, usually obeying power-law. Especially, in the typical model of ultra-slow-roll inflation, $\mathcal{P}_{\mathcal{R}}(k)$ increases as $\propto k^{4}$ before it reaches the maximum Byrnes et al. (2019). There might be modulated oscillations of period $\sim\pi$ if the transition from slow-roll to ultra-slow-roll is abrupt Starobinsky (1992); Biagetti et al. (2018); Cole et al. (2022); Pi and Wang (2023). The UV behavior of the power spectrum depends on the shape of the second slow-roll potential, usually displaying another power-law $\mathcal{P_{R}}\propto k^{n_{s,2}-1}$. If the ultra-slow-roll stage ends sharply, the power spectrum drops off a little bit with new modulated oscillations with a larger period of $\sim\pi e^{N_{\mathrm{USR}}}$ Pi and Wang (2023). According to our discussion in Section III.3, these oscillations mark the sharp end of ultra-slow-roll stage, which are accompanied with the logarithmic relation (30). Based on the analytical expression of $\mathcal{P}_{\mathcal{R}}(k)$ as well as the statistical properties, we can draw the GW spectrum associated with a fixed PBH abundance. Currently, there is no satisfactory method of calculating PBH abundance for such a broad spectrum with non-Gaussianity, so we choose the same parameters as in the narrow peak case (marked “USR” in Table 1), aiming to demonstrate the feature but not the exact amplitude of the induced GW spectrum in the ultra-slow-roll inflation. This is shown in the right panel of Fig. 7. For comparison, we also draw the induced GW from Starobinsky’s linear potential model (with the narrow-peak amplitude), which represents the smooth end of ultra-slow-roll stage with negligible non-Gaussianity. We can see that similar to the $\delta$-function peak case, non-Gaussianity enhances the UV part of the induced GW spectrum, which makes the primary peak of the GW spectrum obscure. A drop off with modulated oscillations can be recognized as a smoking gun for the non-Gaussianity in the ultra-slow-roll inflation shown in (30), which are detectable by DECIGO and BBO when PBHs constitute all the dark matter.

### v3

@@ I Overview
A related phenomenon is the stochastic gravitational wave background induced by the enhanced curvature perturbation. In General Relativity, the curvature perturbation $\mathcal{R}$ and tensor perturbation $h_{ij}$ are only coupled at nonlinear order, and the tensor-scalar-scalar type coupling $\sim h_{ij}\partial_{i}\mathcal{R}\partial_{j}\mathcal{R}$ can induce second-order GWs Matarrese et al. (1993, 1994, 1998); Noh and Hwang (2004); Carbone and Matarrese (2005); Nakamura (2007); Ananda et al. (2007); Osano et al. (2007); Baumann et al. (2007)  [^2] For GWs induced by other couplings like scalar-tensor-tensor or tensor-tensor-tensor, see for instance Refs. Chang et al. (2023); Yu and Wang (2023); Bari et al. (2024); Picard and Malik (2023).. Therefore, the curvature perturbation can source the GWs after it reenters the Hubble horizon, which is called scalar induced secondary GWs. The peak frequency, as well as the PBH mass generated by such a perturbation, are determined by the horizon scale at the reentry moment, which are connected by Saito and Yokoyama (2009)
@@
which is also used to describe the curvature perturbation on small scales for PBH formation Byrnes et al. (2012); Young and Byrnes (2013); Tada and Yokoyama (2015); Young and Byrnes (2015); Young et al. (2016); Franciolini et al. (2018); Ando et al. (2018a); Atal and Germani (2019); Passaglia et al. (2019); Özsoy and Tasinato (2023) as well as the induced GWs Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020). However, in many models, the non-linear parameter $f_{\mathrm{NL}}$ reaches $\mathcal{O}(1)$ Namjoo et al. (2013); Martin et al. (2013); Chen et al. (2013); Motohashi et al. (2015); Davies et al. (2022); Namjoo (2023); Namjoo and Nikbakht (2024), so the higher orders in (2) might be important, which are highly model-dependent. These higher order terms must be taken into account in the PBH formation. Fortunately, by $\delta N$ formalism, a fully nonlinear curvature perturbation can be derived for single field inflation and for the curvaton scenario, of which the PDF of $\mathcal{R}$ can be calculated analytically. Various methods are developed to calculate the PBH abundance for such a PDF, including the Press-Schechter-type formalism Biagetti et al. (2021); Gow et al. (2022); Ferrante et al. (2022) and the theory of peaks Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022). The Press-Schechter-type formalism for Gaussian curvature perturbation is thoroughly reviewed in Chapter 6 Young (2024), which is straightforward to be further extended to non-Gaussian case.
Considering non-Gaussianity in the PBH formation and induced GW generation is important in many theoretical and observational issues in cosmology.
@@ II.2.2 Threshold
We see that depending on whether $\mathscr{C}_{\ell}<4/3$ or $\mathscr{C}_{\ell}>4/3$, the maximum $\mathscr{C}(r_{m})$ corresponds to the maximum or minimum of $\mathscr{C}_{\ell}$. Only the former case, dubbed type I fluctuation Kopp et al. (2011), is consistent with the linear approximation. Type II fluctuations with $\mathscr{C}_{\ell}>4/3$ are usually rarer, which we do not consider in this paper. Given the profile-dependent threshold on $\mathscr{C}$, the corresponding threshold on the linear compaction function $\mathscr{C}_{\ell}$ can be solved by (12)
@@ III.3 A general analysis on single-field inflation
$\mathcal{R}$ $=\delta N_{1}\left(\{\varphi,\pi\}\to\{\varphi_{t},\pi_{t}\}\right)+\delta N_{2}\left(\{\varphi_{t},\pi_{t}\}\to\{\varphi_{e},\pi_{e}\}\right),$ (44)
$\delta N_{1}$ $=\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi+\lambda_{\mp}\delta\varphi}{\pi+\lambda_{\mp}\varphi}\right)-\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda_{\mp}\varphi_{t}}\right),$ (45)
@@
Figure 4: The schematic potential (left panel) and the phase portrait (right panel) of the “smooth transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=-0.01$, $\varphi_{m}=0$. For fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$ (in the unit of $H$), respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Equal-$N$ lines are parallel to the fast-roll streams (i.e. $\pi\propto-3\varphi$) with almost equal intervals, which are not drawn for visual clearance. This clearly shows that the contribution to $\delta N$ from the ultra-slow-roll stage is negligible, and the contribution from the later stage preserves slow-roll-type Gaussianity.
@@
This result is easy to understand by looking at the phase portrait. For the “smooth transition” case, although the potential behaves differently for $\varphi<\varphi_{t}$ and $\varphi>\varphi_{t}$, but the non-attractor solution ($\pi=-3\varphi$) in the flat plateau is almost the same as the non-attractor solution in a slow-roll potential ($\pi=-\lambda_{+}^{\prime}\varphi\approx(3-\eta^{\prime})\varphi$). Therefore, this case is similar to the bumpy potential we studied in Section III.2, and (42) tells us $\mathcal{R}$ is nearly Gaussian with $f_{\mathrm{NL}}=-(6/5)\lambda_{-}^{\prime}$. The phase portrait Fig. 4 looks similar to that of a slow-roll inflation, as the ultra-slow-roll trajectories are similar to the fast-roll trajectories in an ordinary slow-roll inflation, which are parallel to the equal-$N$ lines. The difference in the $e$-folding number for two nearby trajectories are proportional to the difference in $\varphi$, therefore $\delta N\propto\delta\varphi$ which preserves Gaussianity. The well studied Starobinsky model Starobinsky (1992); Leach et al. (2001); Biagetti et al. (2018); Tasinato (2021); Pi and Wang (2023); Domènech et al. (2023) belongs to this limit.
Figure 5: The schematic potential (left panel) and the phase portrait (right panel) of the “sharp transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=0.01$, $\varphi_{m}=100$ (in the unit of $H$). For the fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$, respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Near the origin, the trajectories merge to a unique fast-roll trajectory with different $\pi$, which contribute negligibly to $\delta N$ in the subsequent evolution. The totel $\delta N$ comes mainly from the ultra-slow-roll stage, given by (30).
For the sharp transition case, as is shown in Fig. 5, the inflaton experiences an acceleration after the ultra-slow-roll stage, therefore the nearby trajectories, slightly shifted by the velocity difference $\delta\pi_{t}$ generated during the ultra-slow-roll stage, can also induce difference in $e$-folding number $\delta N$ in the later slow-roll stage. However, as $\pi_{t}\ll 1$, trajectories with different $\pi_{t}$’s near the origin are almost on the same fast-roll trajectory in the slow-roll era for $\varphi>\varphi_{t}$, as the trajectories near the $\varphi$-axis are perpendicular to it. On this trajectory, the difference in $N$ caused by $\delta\pi_{t}$ is negligible, and $\varphi=\varphi_{t}$ is approximately an equal-$N$ surface. Following the discussion in Section III.1, we will arrive at the non-Gaussian relation (30) straightforwardly.
@@ III.4 The curvaton scenario
where $r$ is roughly the cosmological density parameter of the curvaton $\chi$ at its decay. Apparently such a nominal $f_{\mathrm{NL}}$ is not enough to calculate the PBH abundance, as $f_{\mathrm{NL}}$ is always larger than 1.
@@
The fully non-Gaussian curvature perturbation (49) seems complicated, but it has two interesting limiting cases for $r\ll 1$ and $r\to 1$, both of which have fruitful physical implications. In the former limit, the curvaton energy density is negligible at its decay, and (49) can be expanded for small $r$:
@@
$\zeta=\frac{1}{3}\ln A=\frac{2}{3}\ln\left|1+\frac{\delta\chi}{\chi}\right|.$ (54)
@@
$\zeta=\zeta_{\chi}(t,\mathbf{x})=\int^{\rho_{\chi}(t,\mathbf{x})}_{\rho_{\chi}(t)}\frac{\mathrm{d}\widetilde{\rho}_{\chi}}{3\widetilde{\rho}_{\chi}}=\frac{1}{3}\ln\frac{\rho_{\chi}(t,\mathbf{x})}{\rho_{\chi}(t)}=\frac{2}{3}\ln\left|1+\frac{\delta\chi}{\chi}\right|,$ (55)
@@ IV PBH abundance with non-Gaussian curvature perturbation
where $\varrho$ is the correlation between $X$ and $Y$, $j_{0}(z)=\sin z/z$ is the zeroth order spherical Bessel function, and we assume a spherical $\zeta$ profile of
@@
Another solution with $\mathscr{C}_{\ell}>4/3$ is discarded as it is for the Type II fluctuation. With (66), to change the variable from $\mathscr{C}_{\ell}$ to $M$, we have
@@
Apparently for $\lambda\to 0$, it degenerates to the Gaussian PDF. Roughly speaking, the variance of PDF (74) increases for a negative $\lambda$ (i.e. positive nominal $f_{\mathrm{NL}}$), which enhances PBH formation. As $\mathscr{C}_{\ell}$ increases, the PDF (74) finally approaches the Cauchy distribution $\mathbb{P}(\mathscr{C}_{\ell})\propto\mathscr{C}_{\ell}^{-2}$, which, however, is far from the Type-I-PBH formation region of $30.872<\mathscr{C}_{\ell}<4/$ 3 Biagetti et al. (2021). We calculated the PBH mass function $f_{\mathrm{PBH}}(M)$ for such a monochromatic power spectrum (71) in the ultra-slow-roll inflation ($\lambda=-3$), shown in the left panel of Fig. 6. We choose a typical PBH peak mass of $M_{k^{*}}=10^{-12}M_{\odot}$, and normalize the amplitude of the power spectrum $\mathcal{A_{R}}$ to be $4.151077\times 10^{-3}$, such that $f_{\mathrm{PBH}}=1$, i.e. PBHs are all the dark matter. Interestingly, as is clearly seen in Fig. 6, for the narrow peak case, the critical infrared scaling $f_{\mathrm{PBH}}(M)\propto M^{1+1/\gamma}\sim M^{3.78}$ which originates from the universal critical behavior is almost parallel to the constraints from the 21-cm signal which is proportional to $M^{4}$ Mittal et al. (2022). So the entire asteroid-mass window is open for narrow-peak power spectrum even with non-Gaussianity.
@@ V Induced Gravitational Waves
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}\mathrm{g}}(vk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\lambda\int^{1+\lambda}_{|1-\lambda|}\mathrm{d}\rho~{}\frac{\mathcal{P}_{\mathcal{R}\mathrm{g}}(\lambda vk)\mathcal{P}_{\mathcal{R}\mathrm{g}}(\rho vk)}{\lambda^{2}\rho^{2}}+\cdots\right)+\cdots.$ (83)
@@
Figure 7: Left: The spectrum of GWs induced by the enhanced narrow-peak power spectrum (71). The amplitude $\mathcal{A_{R}}$ for different parameters follows Table 1, such that PBHs $\sim 10^{-12}M_{\odot}$ are all the dark matter, shown in Fig.6. From right to left, the orange, magenta dashed, brown dot-dashed, cyan thick, blue dashed, purple dotted, black thick, gray, and black dotted curves are for $f_{\mathrm{NL}}=10^{3}$, $f_{\mathrm{NL}}=10^{2}$, $f_{\mathrm{NL}}=10$, ultra-slow-roll, $f_{\mathrm{NL}}=5/2$, $f_{\mathrm{NL}}=1$, Gaussian case ($f_{\mathrm{NL}}=0$), curvaton scenario with $r\to 1$, and $f_{\mathrm{NL}}=-5/4$, respectively. The curves are drown by Adshead et al. (2023), following Adshead et al. (2021), together with the designed power-law integrated sensitivity curve of space-borne interferometers LISA (gray thick) Bartolo et al. (2016), Taiji (gray dashed) Wang and Han (2021), TianQin (gray dotted) Liang et al. (2021), and DECIGO (gray thin) Kawamura et al. (2006, 2011). Right: Two more realistic signals, the spectrum of GWs induced by ultra-slow-roll inflation with a sharp transition to slow-roll (blue), and a smooth transition to slow-roll, i.e. Starobinsky’s linear potential model (red). The right panel was drawn by (83) with the help of the SIGWfast Witkowski (2022) package.
@@
In the first expression, the numerical coefficient of the Gaussian contribution (“6.4”) comes from Ref. Pi and Sasaki (2020) with smoothing scale $\Delta=0.01$, while the coefficients of the $F_{\mathrm{NL}}^{2}$-term (“3.7”) and $F_{\mathrm{NL}}^{4}$-term (“3.9”) come from Ref. Adshead et al. (2021) (also with a width of $\Delta=0.01$). In the second equality, we use (79) for PBH as all the dark matter. From the above fitting formula, it is easy to see that $\Omega_{\mathrm{GW,peak}}h^{2}$ has a minimum of $\Omega_{\mathrm{GW,peak}}^{\text{(min)}}h^{2}\sim 8.4\times 10^{-11}$ when $f_{\mathrm{NL}}\sim 63$. After that, $\Omega_{\mathrm{GW,peak}}h^{2}$ only increase slightly, and finally approaches a constant of $1.1\times 10^{-10}$. This minimum is still well above the sensitivity curves of space-borne interferometers. Therefore, the detectability of scalar-induced GWs is robust against the nonlinear parameter $f_{\mathrm{NL}}$, $\lambda_{-}$ (see (34) and (42)), or $r$ (see (48) and (49)), which made it an important scientific goal of LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021) to probe the PBH dark matter.
@@
We should warn that the $\delta$-function peak we considered above is not physical. For more realistic models, the enhancement and decay of the power spectrum around the peak are milder, usually obeying power-law. Especially, in the typical model of ultra-slow-roll inflation, $\mathcal{P}_{\mathcal{R}}(k)$ increases as $\propto k^{4}$ before it reaches the maximum Byrnes et al. (2019). There might be modulated oscillations of period $\sim\pi$ if the transition from slow-roll to ultra-slow-roll is abrupt Starobinsky (1992); Biagetti et al. (2018); Cole et al. (2022); Pi and Wang (2023). The UV behavior of the power spectrum depends on the shape of the second slow-roll potential, usually displaying another power-law $\mathcal{P_{R}}\propto k^{n_{s,2}-1}$. If the ultra-slow-roll stage ends sharply, the power spectrum drops off a little bit with new modulated oscillations with a larger period of $\sim\pi e^{N_{\mathrm{USR}}}$ Pi and Wang (2023). According to our discussion in Section III.3, these oscillations mark the sharp end of ultra-slow-roll stage, which are accompanied with the logarithmic relation (30). Based on the analytical expression of $\mathcal{P}_{\mathcal{R}}(k)$ as well as the statistical properties, we can draw the GW spectrum associated with a fixed PBH abundance. Currently, there is no satisfactory method of calculating PBH abundance for such a broad spectrum with non-Gaussianity, so we choose the same parameters as in the narrow peak case (marked “USR” in Table 1), aiming to demonstrate the feature but not the exact amplitude of the induced GW spectrum in the ultra-slow-roll inflation. This is shown in the right panel of Fig. 7. For comparison, we also draw the induced GW from Starobinsky’s linear potential model (with the narrow-peak amplitude), which represents the smooth end of ultra-slow-roll stage with negligible non-Gaussianity. We can see that similar to the $\delta$-function peak case, non-Gaussianity enhances the UV part of the induced GW spectrum, which makes the primary peak of the GW spectrum obscure. A drop off with modulated oscillations can be recognized as a smoking gun for the non-Gaussianity in the ultra-slow-roll inflation shown in (30), which are detectable by DECIGO and BBO when PBHs constitute all the dark matter.